A copper block is removed from a 300 ∘C oven and dropped into 1.30 kg of water at 22.0 ∘C. The water quickly reaches 32.0 ∘C and then remains at that temperature. What is the mass of the copper block? The specific heats of copper and water are 385 J/(kg⋅K) and 4190 J/(kg⋅K) respectively. Express your answer to three significant figures and include the appropriate units.
Answer #1

Specific heat of copper = C1 = 385 J/(kg.K)

Specific heat of water = C2 = 4190 J/(kg.K)

Mass of the copper block = m1

Mass of the water = m2 = 1.3 kg

Initial temperature of the copper block = T1 = 300 oC

Initial temperature of the water = T2 = 22 oC

Final temperature of the system = T3 = 32 oC

The heat lost by the copper block is equal to the heat gained by the water.

m1C1(T1 - T3) = m2C2(T3 - T2)

m1(385)(300 - 32) = (1.3)(4190)(32 - 22)

m1 = 0.528 kg

Mass of the copper block = 0.528 kg